OpenStudy (anonymous):

Find the distance between the points (11, 6) and (0, -2)

OpenStudy (anonymous):

do u know the distance fomula?

OpenStudy (anonymous):

i do not. that would help. ugh... I hate when the teacher doesn't have enough time to go over what is on the homework.

OpenStudy (anonymous):

\[\sqrt{\left( x2 - x1 \right)^{2} + \left( y2 - y1 \right)^{2}}\]

OpenStudy (anonymous):

if the point s are (x1,x2) and (y1,y2)

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

You're Welcome

OpenStudy (anonymous):

\[\text{The distance between two points }(x_{1},y_{1})\;\;and\;\;(x_{2},y_{2})\text{ is } \sqrt{(x_{1}-x_{2})^{2}+(y_{1}-y_{2})^{2}}.\]